Consider two genes that are independent, the A gene and the B gene. Each gene has two alleles (A gene: A1 and A2, B gene: B1 and B2). The frequency of A1 allele at the A gene is 0.67 and the frequency of the B2 allele at the B gene is 0.21. What is the probability of an individual in a randomly mating population of having A1A2B1B2 as their genotype?

(calculate to 4 decimal places
Answer #1

since probability of having A1A2 allele =2*pq =2*0.67*0.33=0.4422

and probability of having B1B2 allele=2*0.21*0.79 = 0.3318

probability of an individual in a randomly mating population of having A1A2B1B2 as their genotype

=P(A1A2)*P(B1B2) =0.4422*0.3318 =0.1467